36-303: Sampling, Surveys and Society Notes on Team Working Assignments Tue Feb 9, 2010

Here is a revised schedule of Team Project Assignments, up through Spring Break (I have been falling a little behind anyway, and the snow days make it even more important that we be efficient about moving forward).

So far…

I.0 Teams Formed (Tue Jan 19)

I.1 Propose Two Topics (Tue Jan 26)

I.2 Revise proposals (Thu Feb 4)

I.3 Choose Topic (Thu Feb 11: one short email per team to brian@stat.cmu.edu) [Do not worry about IRB application or draft questions yet]

Upcoming…

The remaining Team Project Assignments will more or less follow the shorter “designing a sample survey” outline from the beginning of the semester. DATES AND DETAILS HAVE CHANGED; PLEASE READ BELOW.

I.4 Target Pop, Sampling Plan, Sampling Frame, Nonresponse Plan (Tue Feb 16; \underline{one page}, emailed to brian@stat.cmu.edu)

• Clearly define a \underline{target population} that you can reach in your survey

• Clearly discuss your \underline{sampling plan}: How will you obtain a representative sample from the the target population you defined? This should involve:

– Clearly define a \underline{sampling frame}: a list of persons in the target population from which you can draw a random sample. Acceptable frames include: (a) list of email addresses from which you can draw an SRS without replacement; (b) set of phone numbers or prefixes that you can use for random-digit dialling (this is again an SRS without replacement); (c) list or map of blocks, addresses, or similar geographic information you can use to randomly select houses or businesses to approach; etc. There are very few cases where a non-random sample with no clearly defined frame (e.g. man-on-the-street interviews) is acceptable. For man-on-the-street interviews, careful design of times/dates/locations, and method of approach people “on the street” (or on-campus, etc.) will be needed.

– Clearly discuss how you will use the sampling frame to obtain a representative sample.

– Clearly discuss how you will deal with \underline{nonresponse}:

∗ What will you do about reminding or following-up nonresponders?

∗ If you suspect your sample is not representative of the target population, what adjustments do you anticipate making?

Team Working Agreements (Tue Feb 23, in email to brian@stat.cmu.edu and on paper.) [see separate handout]

II.5a Sampling Scheme & Question Design (Thu Feb 25; email attachment to brian@stat.cmu.edu; I will only provide feedback if needed)

• Items K, L, M on the shorter “designing a sample survey” handout

• Items K and M should be 1/2 page each. Item L should be however long is needed.

• For item M, please also include information on oversampling to compensate for nonresponse. For example, if you anticipate a 25% response rate and you want a sample size of 120, how many respondents should you approach?

II.5b Pretest & Revise Questions (Tue Mar 2; email brian@stat.cmu.edu ONLY part O. I will provide feedback only if needed.)

• Items N, O on the shorter “designing a sample survey” handout

II.6 Final IRB & Project Plan (Thu Mar 4; email attachments to brian@stat.cmu.edu);

• Final, full project proposal (items A-M on the shorter “designing a sample survey”

handout).

– From this proposal, anyone outside our class should be able to read and understand

completely what you are proposing to do.

• IRB form filled out completely

– You can get a copy of this form from the course website or from

http://www.cmu.edu/osp/regulatory-compliance/human-subjects.html

– To save time, you do NOT need to do the IRB Training Certificate indicated in blocks 2 and 3 of the IRB form. Ignore those checkboxes (but do fill out the rest of those blocks).

• A draft informed consent statement for your survey.

Turn in all three to me, not CMU IRB.